OpenStudy (anonymous):

Help me find the x-intercept of the equation y=(x+2)^3 + 5, show working & give exact value thanks. I know to find the x-int you have to make y=0 but I think I've got the wrong answer.

OpenStudy (anonymous):

Hi JayDS, when you put y=0 the equation becomes like (x+2)^3+5=0. Therefore it is a three degree polynomial with 3 roots may be real or complex. so, solving that equation, the roots u get are, x1 = -3.71, x2 = -1.145 + 1.4809*i and x3 = -1.145 - 1.4809*i.

OpenStudy (anonymous):

Ok, I kind of get it but can you explain the real or complex part and how you got the values of x1, x2 and x3?

OpenStudy (anonymous):

It is not possible to use the trial and error method to get the values i mean the roots for that polynomial.. so i used a online calculator to get the roots

OpenStudy (anonymous):

kk. thanks.

OpenStudy (anonymous):

check this link to get the formulae for the roots of 3rd degree polynomial ... http://www.mathportal.org/formulas/algebra/solalgebric.php